OpenStudy (anonymous):

The number N of cars produced at a certain factory in 1 day after t hours of operation is given by N(t) = 100t – 5t2, 0 ≤ t ≤ 10. If the cost C (in dollars) of producing N cars is C(N) = 15,000 + 8000N, find the cost C as a function of the time t of operation of the factory.

OpenStudy (anonymous):

\[C(N(t))=C(100t-5t^2)=15,000+8000(100t-5t^2)\]